OpenStudy (dellgato):

Find the perimeter of a rectangle with side lengths (2x + 12) ft. and (4x - 13) ft.

TheSmartOne (thesmartone):

Perimeter of a rectangle = 2(Length + Width)

OpenStudy (dellgato):

ok

OpenStudy (dellgato):

check my answer in a min

TheSmartOne (thesmartone):

sure :)

OpenStudy (dellgato):

So would it be (4x+24) and (8x-13) or (12x-37)

TheSmartOne (thesmartone):

You need to first add the length and width 2x + 12 + 4x - 13 2x + 4x = ? 12 - 13 = ? And then you m ultiply the whole thing by 2. You'll have to distribute the 2. Remember the distributive property? a(b - c) = ab - ac

OpenStudy (dellgato):

Ok so 2x+4x is 6x and 12-13 is -1 so what after that?

OpenStudy (dellgato):

TheSmartOne

jimthompson5910 (jim_thompson5910):

@Dellgato you need to stick an @ sign before the user name to get his attention like this @TheSmartOne

jimthompson5910 (jim_thompson5910):

Though he appears to be offline now

OpenStudy (dellgato):

Ok, Thank you!

jimthompson5910 (jim_thompson5910):

once you get to 6x-1, you would distribute the 2 through 2*(6x-1) = ???

OpenStudy (dellgato):

(12 x - 2)

OpenStudy (leenathan):

...

jimthompson5910 (jim_thompson5910):

12x-2 is the perimeter

OpenStudy (dellgato):

Ok, thanks a lot

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (leenathan):

please close this question now

OpenStudy (dellgato):

How?

OpenStudy (leenathan):

go to the top of your qustion and you will see that button you should do this after your don with a qustion

TheSmartOne (thesmartone):

Great work! And thanks, Jim! :)